OpenStudy (anonymous):

A bacteria culture begins with 13 bacteria which triple in amount at the end of every hour. How many bacteria exist in the culture at the end of the 10th hour?

OpenStudy (anonymous):

13*3^10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

dy/dx = 13+3x y = 13x + 3x^2/2 after 10 hours 13(10) + 3(10)^2/2 = 280 bacteria

OpenStudy (anonymous):

Thanks but thats not an answer choice

OpenStudy (anonymous):

my answer is wrong :)